OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95.

OpenStudy (anonymous):

a 203 b 197 c 168 d 163

OpenStudy (anonymous):

197

hartnn (hartnn):

the general 'n' term is given by \(\huge a_n=a_1+(n-1)d\) so a18=-7+17d=95 can u find 17 d from here? now a35=-7+34d=-7+2(17d)=....

OpenStudy (anonymous):

oh okay so would it be 197 then?

hartnn (hartnn):

yup, thats correct :)

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome :)